OpenStudy (anonymous):

The football team has played 12 games so far this season. They have won 8 games. What percent of their games has the team lost? 33 1/3% 67% 75% 80%

OpenStudy (anonymous):

33 1/3 %

OpenStudy (anonymous):

the games lost = 12 -8 = 4 x% of 12 = 4 x/100 * 12 = 4 x/100 = 4/12 = 1/3 x = 1/3 * 100 = 100/3 = 33.33 % or 33 1/3 %

OpenStudy (anonymous):

good luck...now me off to sleep :)